# SPHEREx Pre-Perihelion Mapping of $\mathrm{H_{2}O}$, $\mathrm{CO_{2}}$, and $\mathrm{CO}$ in Interstellar Object 3I/ATLAS

### v1

@@ Abstract
From 01- to 15-Aug-2025UT, the SPHEREx spacecraft observed interstellar object 3I/ATLAS. Using $R=40\mathrm{-}130$ spectrophotometry at $\lambda=0.7\mathrm{-}5\,\mathrm{\mu m}$, light curves, spectra, and imaging of 3I were obtained. From these, robust detections of water gas emission at $2.7\mathrm{-}2.8\,\mathrm{\mu m}$ and $\mathrm{CO}_{2}$ gas at $4.23\mathrm{-}4.27\,\mathrm{\mu m}$ plus tentative detections of $^{13}\mathrm{CO}_{2}$ and $\mathrm{CO}$ gas were found. A slightly extended $\mathrm{H_{2}O}$ coma was detected, and a huge $\mathrm{CO}_{2}$ atmosphere of extending out to at least $4.2\times 10^{5}\,\mathrm{km}$ was discovered. Gas production rates for $\mathrm{H_{2}O}$, $^{12}\mathrm{CO}_{2}$, $^{13}\mathrm{CO}_{2}$, and $\mathrm{CO}$ were $Q_{gas}=3.2\times 10^{26}\pm 20\,\mathrm{\%}$, $1.6\times 10^{27}\pm 10\,\mathrm{\%}$, $1.3\times 10^{25}\pm 25\,\mathrm{\%}$, and $1.0\times 10^{26}\pm 25\,\mathrm{\%}$, respectively. Co-addition of all $\lambda=1.0\mathrm{-}1.5\,\mathrm{\mu m}$ scattered light continuum images from produced a high SNR image consistent with an unresolved source. The scattered light lightcurve showed $\lesssim 15\,\mathrm{\%}$ variability over the observation period. The absolute brightness of 3I at $1.0\mathrm{-}1.5\,\mathrm{\mu m}$ is consistent with a $2.5\,\mathrm{km}$ radius nucleus surrounded by a 100 times brighter coma. The $1.5\mathrm{-}4.0\,\mathrm{\mu m}$ continuum structure shows a strong feature commensurate with water ice absorption seen in KBOs and distant comets. The observed cometary behavior of 3I, including its preponderance of $\mathrm{CO}_{2}$ emission, lack of $\mathrm{CO}$ output, small size, and predominance of large icy chunks of material in a flux-dominant coma is reminiscent of the behavior of short period comet 103P/Hartley 2, target of the NASA Deep Impact extended mission in 2010 and a “hyperactive comet” near the end of its outgassing lifetime. This correspondence places 3I closer to barely- or non-active 1I/Oumuamua than primitive, ice rich 2I/Borisov, suggesting that ISOs are often highly thermally processed before ejection into the ISM.
@@ 4 Results
4.1 Lightcurves
Lightcurve analysis of 3I’s temporal behavior (Fig. 1) shows very little variability of $\lesssim 15\,\mathrm{\%}$ from hours to days timescales, which is also confirmed by independent SOLO observations. This upper limit of variation is also consistent with previous reports (Seligman et al., 2025; Kareta et al., 2025; Beniyama, 2025). The fortuitous lack of temporal variability, explained by an object with a small, likely dark nucleus surrounded by a very bright and high surface area coma (see Section 4.6 below), meant that we could quickly cull out any background contaminated images and proceed straight to analysis of the data.
4.2 Spectra
@@ 4.3 Gas Production Rates
[…]
@@ 4.4 Images
Images at different wavelength ranges probe the distribution of dust and important spectral features by molecules (Sect. 4.2). The large extension seen in $^{12}\mathrm{CO}_{2}$ versus $\mathrm{H_{2}O}$ is not surprising given the orders of magnitude higher measured flux emission from $\mathrm{CO}_{2}$ gas, the object’s location well within the $\sim 12\,\mathrm{au}$ $\mathrm{CO}_{2}$ ice line of the solar system, and the long lifetime ($\sim 6$ days) vs photolysis and SWCX compared to water ($\sim 1$ day). However, the falloff of the $\mathrm{CO}_{2}$ coma with respect to projected radial distance $\rho$ is definitely faster/steeper than $1/\rho$, implying that there is either acceleration of the outflow occurring or there is destruction of the outflowing $\mathrm{CO}_{2}$ going on.
@@ 5 Discussion
Figure 4: (a) $\mathrm{CO}/\mathrm{CO}_{2}$ gas production rate ratios of solar system comets (Harrington Pinto et al., 2022). The $\mathrm{CO}$ or $\mathrm{CO}_{2}$ dominance criteria of $1\pm 0.33$ is indicated as horizontal shadow (Harrington Pinto et al., 2022). The ratio of 3I found by SPHERExis $\lesssim 0.013$, making it extremely $\mathrm{CO}$ poor and equivalent to the most thermally processed comet in the 2022 survey, 103P. (b) Pre-perihelion optical lightcurves observed for comet 103P during the 2010–2011 apparition and 3Iin May - Oct 2025. Published data from (Jewitt & Luu, 2025; Tonry et al., 2025; Ye et al., 2025) have been shifted to match the absolute magnitude of Minor Planet Center (MPC) Database [^7] https://www.minorplanetcenter.net/db_search/show_object?object_id=3I for 3I and https://www.minorplanetcenter.net/db_search/show_object?object_id=103P for 103P. , the T-mag from JPL Horizons ephemerides service [^8] https://ssd.jpl.nasa.gov/horizons/app.html, and the data kindly shared by S. Yoshida (2025999http://www.aerith.net/comet/catalog/0003I/2025N1.html, priv. comm.). There is very good agreement between the two inbound lightcurve trends. An inherent spread of $\sim 0.5\,\mathrm{mag}$ in the measurements between nights and between observers can be seen. (c) Close flyby imaging of the $\sim 1.5\times 0.6\times 0.6\,\mathrm{km}$ comet 103P nucleus as obtained by the Deep Impact Extended mission (A’Hearn et al., 2011). A marked bilobate nuclear morphology was seen, with unusually large ($m$ to dm) chunks of $\mathrm{CO}_{2}$ rich ice and gas emanating from the end of the smaller lobe and the middle of the larger lobe. The surrounding coma, filled with large long lived icy grains, was found to be at least 5 times brighter and more actively emitting gas than the nucleus. While the nucleus of 3I has never been resolved, given its similar behavior, it could possibly share similar morphological characteristics. (d) Schematic of the 3I object derived from SPHEREx $\mathrm{CO}_{2}$ photometric mapping. The $2.5\,\mathrm{km}$ radius nucleus (Jewitt et al., 2025) was not resolved at the SPHEREx pixel scale of $6.15\arcsec\approx 11,600\,\mathrm{km}$ at $2.6\,\mathrm{au}$ distance, but a notional nucleus image has been included at the origin to give the reader its location and a sense of scale. Two distances have been marked off: The $\sim 38,000\,\mathrm{km}$ (white dashed circle) inside of which we think coma ice is actively sourcing $\mathrm{CO}_{2}$ by sublimation; and $\sim 340,000\,\mathrm{km}$ (orange dashed oval), the minimum radial distance to which we detect $\mathrm{CO}_{2}$ gas at 2-$\sigma$ level.
@@ 5.1 Other Hypotheses
5.2 Home System Implications
@@ 5.3 ISO Evolution
6 Conclusions and Future Work
@@
This paper relies on data obtained by the SPHEREx Observatory (operated by Caltech and JPL on behalf of the National Aeronautics Space Administration under contract 80GSFC18C0011) that is hosted by IPAC as part of the IRSA archive (10.26131/IRSA652 SPHEREx Quick Release Spectral Images - QR2). We are indebted to the SPHEREx project for obtaining additional pointings at 3I, approximately 100 percent more, than their nominal survey would have obtained less than 3 months after the start of the survey by testing and implementing a new commanding protocol. Part of the research described in this paper was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The authors also acknowledge the Texas Advanced Computing Center (TACC)  [^10] http://www.tacc.utexas.edu at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper. {contribution}All listed authors contributed importantly to the creation and production of this manuscript.

### v2

@@ Abstract
[…]
@@ IV Results
IV.1 Lightcurves
Lightcurve analysis of 3I’s temporal behavior (Fig. 1) shows very little variability of $\lesssim 15\,\mathrm{\%}$ from hours to days timescales, which is also confirmed by independent SOLO observations. This upper limit of variation is also consistent with previous reports (Seligman et al., 2025; Kareta et al., 2025; Beniyama, 2025). The fortuitous lack of temporal variability, explained by an object with a small, likely dark nucleus surrounded by a very bright and high surface area coma (see Section IV.6 below), meant that we could quickly cull out any background contaminated images and proceed straight to analysis of the data.
IV.2 Spectra
@@ IV.3 Gas Production Rates
where $N_{\mathrm{gas}}$ is the average column density of the gas within a circular aperture of projected radius $r$ on the sky, centered on the nucleus, and where $v_{\mathrm{gas}}$ is the gas emission velocity. We use $v_{\mathrm{gas}}=(0.85\,\mathrm{km/s})/\sqrt{r_{\mathrm{hel}}\text{ in au}}=0.48\,\mathrm{km/s}$ (Cochran et al., 2012) and $r=11,600$ km here. The average column density is given by (cf. e.g. Konno & Wyckoff, 1989)
@@ IV.4 Images
Images at different wavelength ranges probe the distribution of dust and important spectral features by molecules (Sect. IV.2). The large extension seen in $^{12}\mathrm{CO}_{2}$ versus $\mathrm{H_{2}O}$ is not surprising given the orders of magnitude higher measured flux emission from $\mathrm{CO}_{2}$ gas, the object’s location well within the $\sim 12\,\mathrm{au}$ $\mathrm{CO}_{2}$ ice line of the solar system, and the long lifetime ($\sim 6$ days) vs photolysis and SWCX compared to water ($\sim 1$ day). However, the falloff of the $\mathrm{CO}_{2}$ coma with respect to projected radial distance $\rho$ is definitely faster/steeper than $1/\rho$, implying that there is either acceleration of the outflow occurring or there is destruction of the outflowing $\mathrm{CO}_{2}$ going on.
@@ V Discussion
Figure 4: (a) $\mathrm{CO}/\mathrm{CO}_{2}$ gas production rate ratios of solar system comets (Harrington Pinto et al., 2022). The $\mathrm{CO}$ or $\mathrm{CO}_{2}$ dominance criteria of $1\pm 0.33$ is indicated as horizontal shadow (Harrington Pinto et al., 2022). The ratio of 3I makes it extremely $\mathrm{CO}$ poor and equivalent to the most thermally processed comet in the 2022 survey, 103P. (b) Pre-perihelion optical lightcurves observed for comet 103P during the 2010–2011 apparition and 3Iin May - Oct 2025. Published data from (Jewitt & Luu, 2025; Tonry et al., 2025; Ye et al., 2025) have been shifted to match the absolute magnitude of Minor Planet Center (MPC) Database [^7] https://www.minorplanetcenter.net/db_search/show_object?object_id=3I for 3I and https://www.minorplanetcenter.net/db_search/show_object?object_id=103P for 103P. , the T-mag from JPL Horizons ephemerides service [^8] https://ssd.jpl.nasa.gov/horizons/app.html, and the data kindly shared by S. Yoshida (2025999http://www.aerith.net/comet/catalog/0003I/2025N1.html, priv. comm.). There is very good agreement between the two inbound lightcurve trends. An inherent spread of $\sim 0.5\,\mathrm{mag}$ in the measurements between nights and between observers can be seen. (c) Close flyby imaging of the $\sim 1.5\times 0.6\times 0.6\,\mathrm{km}$ comet 103P nucleus as obtained by the Deep Impact Extended mission (A’Hearn et al., 2011). A marked bilobate nuclear morphology was seen, with unusually large ($m$ to dm) chunks of $\mathrm{CO}_{2}$ rich ice and gas emanating from the end of the smaller lobe and the middle of the larger lobe. The surrounding coma, filled with large long lived icy grains, was found to be at least 5 times brighter and more actively emitting gas than the nucleus. While the nucleus of 3I has never been resolved, given its similar behavior, it could possibly share similar morphological characteristics. (d) Schematic of the 3I object derived from SPHEREx $\mathrm{CO}_{2}$ photometric mapping. The $2.5\,\mathrm{km}$ radius nucleus (Jewitt et al., 2025) was not resolved at the SPHEREx pixel scale of $6.15\arcsec\approx 11,600\,\mathrm{km}$ at $2.6\,\mathrm{au}$ distance, but a notional nucleus image has been included at the origin to give the reader its location and a sense of scale. Two distances have been marked off: The $\sim 38,000\,\mathrm{km}$ (white dashed circle) inside of which we think coma ice is actively sourcing $\mathrm{CO}_{2}$ by sublimation; and $\sim 340,000\,\mathrm{km}$ (orange dashed oval), the minimum radial distance to which we detect $\mathrm{CO}_{2}$ gas at 2-$\sigma$ level.
@@ V.1 Other Hypotheses
V.2 Home System Implications
@@ V.3 ISO Evolution
VI Conclusions and Future Work
@@
This paper relies on data obtained by the SPHEREx Observatory (operated by Caltech and JPL on behalf of the National Aeronautics Space Administration under contract 80GSFC18C0011) that is hosted by IPAC as part of the IRSA archive (10.26131/IRSA652 SPHEREx Quick Release Spectral Images - QR2). We are indebted to the SPHEREx project for obtaining additional pointings at 3I, approximately 100 percent more, than their nominal survey would have obtained less than 3 months after the start of the survey by testing and implementing a new commanding protocol. Part of the research described in this paper was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). The authors also acknowledge the Texas Advanced Computing Center (TACC)  [^10] http://www.tacc.utexas.edu at The University of Texas at Austin for providing computational resources that have contributed to the research results reported within this paper.
{contribution}
All listed authors contributed importantly to the creation and production of this manuscript.